# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

### v1

@@ 5.4.3 Accuracy-interpretability trade-off
Optimal ($\operatorname{max-depth}=4$)
Data set $n$ $p$ Binom. Accuracy CART
Space_ga 3107 6 68.7 68.7 67.5
Phoneme 3172 5 69.8 69.7 68.4
Pollen 3848 5 50.6 50.6 50.5
Analcatdata_supreme 4052 7 91.0 90.9 90.8
Abalone 4177 8 69.4 69.4 68.7
Wilt 4839 5 96.4 96.4 96.2
Banana 5300 2 69.8 69.8 69.5
Delta_ailerons 7129 5 81.8 81.7 81.7
Bank8fm 8192 8 80.1 80.0 79.7
Puma8nh 8192 8 70.2 70.2 69.5
Visualizing_soil 8641 4 84.2 84.2 84.2
Delta_elevators 9517 6 75.2 75.2 74.8
Bank-marketing 10578 7 68.2 68.2 67.4
Mozilla4 15545 5 85.8 85.8 85.7
Pulsar-dataset-htru2 17898 8 95.7 95.7 95.7
California 20634 8 70.5 70.5 68.9
California-housing-classi 20640 8 69.8 69.9 68.7
House_8l 22784 8 80.0 80.0 79.3
Bitcoinheist_ransomware 24780 7 66.7 66.7 66.3
Airlines 26969 7 60.6 60.6 60.3
Electricity 38474 8 68.3 68.3 67.4
Click_prediction_small 39926 8 83.2 83.2 83.2
Diabetes130us 71090 7 57.2 57.2 57.0
Run_or_walk_information 88588 6 79.8 79.8 79.4
Wins ($n\leq 250$) 41 28 24
Wins ($2500250<n<$ 2500) 43 34 8
Wins ($n\geq 2500$) 23 20 3
Average rank 1.67 1.79 2.54
Table 3: Size-weighted average up to 16 nodes (SWA16) for optimal and CART. $n$ is the number of samples and $p$ the number of features (before binarization). The results on the data sets with $n\geq 2500$ are shown in detail. The wins and average rank are for all 165 OpenML data sets in our benchmark.
@@
Optimal methods have a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
To test Claim 2, we compute the size-weighted accuracy (SWA) introduced in Section 5.3 to measure the surface under the accuracy-interpretability Pareto front. Table 3 shows SWA up to 16 nodes for ODTs and CART for all real data sets. ODTs are trained with a maximum depth of four and CART is trained without a depth limit. Since Section 3.4 showed that the pessimistic binomial objective (Binom.) found smaller trees with almost the same accuracy as the accuracy objective, we include it as well in the table. For brevity, we show only the results for the largest data sets (with 2500 or more instances) and summarize the results of the smaller data sets by reporting the number of wins.
These results show that on average the optimal approach achieves a higher SWA than CART. The results also repeat the conclusion of Section 3.4, that training optimal methods with objectives other than accuracy can be advantageous. For small data sets the variance is higher and the wins are more evenly distributed, but for larger data sets, the optimal methods significantly outperform CART, thus verifying Claim 2. For large data sets, the training accuracy is close to the test accuracy for highly regularized models which means that optimal decision trees reliably improve over greedy trees.
@@ 5.4.4 Data Efficiency
Figure 16: Results on the synthetic data sets for increasing (a) number of training samples, and (b) number of numeric features.
Fig. 16 shows how ODTs compare with CART for an increasing number of training instances on synthetic data generated from true trees of depth three. For less than 1000 instances, the ODTs are more accurate than both CART with and without a maximum depth limit. For more than 1000 instances, both ODT and CART obtain 100% test accuracy. CART, however, uses 11 leaf nodes to achieve this result, whereas the optimal approach only requires eight (equal to the true tree’s complexity). Both approaches have approximately the same true discovery rate. However, ODT’s false discovery rate is lower. More instances help the ODT method to reduce its false discovery rate.
@@
Figure 18: Actual number of leaf nodes when training CART with depth limit 20 for increasing training samples with and without tuning. These numbers far exceed the eight leaf nodes for Optimal in Fig. 17.
Fig. 17 shows the out-of-sample accuracies for increasing training sample sizes on three large real data sets. We train ODTs with a depth limit of 3, CART with a depth limit of 3, and CART without a depth limit. When the methods are not tuned, the optimal approach overfits on small data sets, obtaining a lower test accuracy than depth-limited CART. However, with tuning, this effect disappears and the ODTs’ accuracy is consistently higher than CART’s (depth limited). The difference in performance between tuning and not tuning diminishes for larger training sets. Without a depth limit, CART continues to increase its accuracy for more data. More data does not help depth-limited CART since, at some point, the greedy decisions on what feature and threshold to split on do not change anymore. In those cases, the added data does not lead to different greedy decisions but only makes them more certain.
However, Fig. 18 also shows that CART continues to grow larger trees with up to tens of thousands of nodes. For the Electricity data set, for example, CART and Optimal have roughly similar accuracy for ten thousand training samples. However, the ODT has eight leaf nodes, whereas CART has over a hundred. These results contradict the observation by Oates and Jensen (1997) that greedy tree methods do not perform much better for more data but do yield larger trees with more data. For these data sets, we observe CART performing much better for more data while also resulting in larger trees.
In conclusion, these results refute Claim 3 that the difference between optimal methods and greedy diminished for more data. CART (without a depth limit) can improve performance over ODT with sufficient data. However, unconstrained CART uses unnecessary splits and can result in trees that are orders of magnitude larger than ODTs. Depth-constrained CART may fail to recover an accurate tree even with large training sets and the difference with ODTs does not diminish. Therefore for both depth-constrained and unconstrained CART, we find that they remain different from ODTs with more data.
@@ 5.4.5 Model Complexity and Training Data
Figure 19: Results on the synthetic data sets for ground truth trees of increasing complexity and training samples $n=50\cdot 2^{d}$, with $d$ the depth of the ground truth tree.
@@ 5.4.6 Overfitting
Figure 20: Results on the synthetic data for increasing (a) feature noise, and (b) class noise.
We addressed overfitting before when observing the results in Fig. 17. These results showed that without hyperparameter tuning, ODTs are more prone to overfitting than greedy approaches when data is sparse. However, with tuning and with the same size constraint, we observe that ODTs perform better than greedy trees on average.
We further analyze the risk of overfitting by comparing ODTs with greedy on the synthetic data with noise. Fig. 20 shows how both approaches respond to increasing feature and class noise. CART has a similar TDR for increasing feature noise as ODT but a much higher FDR throughout all levels of feature noise. CART with a depth constraint has a lower TDR and accuracy. When feature noise is added, it has a sudden jump in tree size, which corresponds with an increase in the FDR. Each method is only slightly affected by the feature noise.
The impact of class noise is stronger, and we see a steep decrease in accuracy. Depth-limited CART again has a sudden jump in the tree size with a small amount of class noise. As more class noise is added, all methods yield smaller trees and observe a reduced TDR. Only the ODTs observe a growth in the FDR.
These results show that ODTs are not more sensitive to noise than greedy trees. Combined with the previous result on learning trees with a small training sample, this proves that Claim 5 is false when ODTs are properly tuned. At a fixed size, ODTs can overfit more than greedy trees but achieve better performance than greedy at a smaller size. Therefore, the regularization effect of a fixed size on optimal and greedy trees is different and cannot be directly compared.
5.4.7 Expected Tree Depth
@@
The expected depth of greedy trees remains (in practice) close to the optimal tree depth, even for complex (large) true tree size (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
Figure 21: The expected depth on the synthetic data is almost always the same.
The expected depth (or question length) is the expected number of branching nodes visited by a test instance. We evaluate Claim 6 about the expected depth of optimal and greedy trees on the synthetic data. Because of the large runtime, we evaluated the synthetic data for true trees of depth 5 on 250 runs rather than the usual 500.
Fig. 21 shows that CART’s expected depth in all cases, for varying training data size, number of features, feature noise, class noise, and ground truth complexity, remains the same as the optimal approach. When CART is depth-constrained, it remains similar to the optimal approach, except when feature noise or some class noise (but not too much) is added. Then, depth-constrained CART yields larger expected depths. Therefore, our results support Claim 6. Although the maximum depth of greedy methods can be significantly larger than the true depth, most deep paths will only be reached by a few samples.
@@ 5.5 Scalability of Optimal Decision Trees
Table 4: Approximate magnitudes of runtimes for training ODTs using STreeD on synthetic data. s for (sub)seconds, m for minutes, and dashes for runtimes over 2 hours.
Since scalability is one of the limitations of ODTs, we test for what problem sizes the use of ODTs is practically feasible. Table 4 provides an overview of runtimes for the optimal method STreeD in terms of seconds, minutes, and hours. We find that training ODTs up to depth four remains practically feasible for data sets up to approximately 250 binary features for 100,000 instances and 150 binary features for one million instances.
Given such runtime results, and the observation that optimal decision trees are 1) specifically good for training small trees up to depth four (Claim 1), 2) for obtaining trees with a good accuracy-interpretability trade-off (Claim 2), and 3) increasingly so for data sets with many instances, we see the best application of ODTs in finding small interpretable trees for medium to large data sets with relatively few but meaningful and well-prepared features.
5.6 Discussion
We have verified and refuted several claims from the literature about optimal and greedy decision tree methods. Optimal methods outperform greedy approaches when constrained with the same depth limit (up to depth four), verifying Claim 1. They obtain solutions with a better accuracy-interpretability trade-off (verifying Claim 2). Though the greedy trees are larger, the average question length for both methods is roughly the same (verifying Claim 6).
However, size-constrained greedy trees remain worse than ODTs with more data, and unconstrained greedy trees need larger trees for similar ODT performance (refuting Claim 3). For increasing ground truth complexity, the greedy approach requires more data than the optimal approach to retain the same accuracy (refuting Claim 4). Though the variance of the accuracy for small data sets is high, this is true for both methods, and we did not observe more overfitting for the optimal approach (refuting Claim 5). These results solidify the value of optimal trees when size-limited trees are desired.
@@ 6 Conclusion
We addressed how to train optimal decision trees (ODTs) and how they compare to greedy approaches and provide recommendations for both (Recommendations 1 and 2).
@@
Since the design of greedy top-down induction (TDI) methods prevents direct accuracy optimization, the literature shows a large variety of learning objectives. We identified and analyzed nine such existing decision tree learning objectives and observed that the concavity of these objectives leads to counter-intuitive results when trained to optimality. Based on this analysis, we propose four novel non-concave objective functions. In contrast to greedy trees, we found that ODTs can optimize out-of-sample accuracy by directly maximizing the in-sample accuracy (provided the data set is large enough), leading to accuracy improvements of more than 1% over objectives such as entropy and Gini impurity. For small data sets, we show that the novel non-concave objectives outperform directly optimizing accuracy because they penalize leaf nodes with close to uniform class distributions more strongly.
@@
We further evaluated four existing and three novel methods for tuning a tree’s complexity across 164 datasets with trees up to depth four, and 131 datasets with trees up to depth five. The experiments show that tuning has a statistically significant impact on the accuracy of at least 1%. The differences between the seven methods are small, specifically for small and large data sets. For medium-sized data sets, tuning the number of nodes, the complexity cost, or the smoothing parameter is best.
@@
Another important contribution of this work is establishing a set of best practices and the supporting analysis framework for comparing ODT and greedy methods. For this, we analyzed previous comparisons and then used the newly established procedure to test six claims from the literature. Our results confirm that (1) ODTs, on average, outperform greedy trees by 1-2% in out-of-sample accuracy for trees trained up to depth four. (2) ODTs, on average, have a better accuracy-interpretability trade-off, which we evaluate with a new metric: the size-weighted accuracy. We refute the claim that (3) differences between greedy and optimal diminish for more data. Depth-constrained greedy methods may fail to recover the true tree, even for large data sets, where ODTs succeed. Unlike previously claimed, (4) greedy methods do not maintain a similar accuracy when the data set size increases linearly with the true tree complexity. ODTs, on the other hand, do. When the complexity of a tree is properly tuned, we find no support for the claim that (5) ODTs are more prone to overfitting than greedy trees. Finally, we find supporting evidence for the claim that (6) the expected depth of greedy trees remains close to the optimal tree.
These results increase our understanding of how to train ODTs and solidify their importance, specifically to obtain small interpretable trees with high accuracy performance.

### v2

@@ 5.4.3 Accuracy-interpretability trade-off
Optimal methods obtain a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
To test Claim 2, we compute the size-weighted accuracy (SWA) introduced in Section 5.3 to measure the surface under the accuracy-interpretability Pareto front. We train ODTs with the top four non-concave and top two concave objectives from Section 3 and CART with the traditional Gini impurity objective. ODTs are trained with a maximum depth of four and CART is trained without a depth limit.
Figure 13: Comparing the test size-weighted average (SWA) of several ODT objectives and CART (see Section 5.3). (a) The non-concave objectives are blue, the concave objectives are orange, and CART is green. (b) The ODT minimum error, accuracy, and M-loss are significantly better than the other approaches. All ODT approaches are significantly better than CART, confirming Claim 2.
Fig. 13 shows that on average the optimal approach achieves a higher SWA than CART, thus verifying Claim 2. For large data sets, the training accuracy is close to the test accuracy for highly regularized models which means that optimal decision trees reliably improve over greedy trees. This result also repeats the conclusion of Section 3.4, that non-strictly-concave objectives are to be preferred over concave objectives for optimal methods.
@@ 5.4.4 Data Efficiency
Fig. 14 shows how ODTs compare with CART for an increasing number of training instances on synthetic data generated from ground-truth trees of depth three. For less than 1000 instances, the ODTs are more accurate than both CART with and without a maximum depth limit. For more than 1000 instances, both ODT and CART obtain 100% test accuracy. CART, however, uses 11 leaf nodes to achieve this result, whereas the optimal approach only requires eight (equal to the true tree’s complexity). Both approaches have approximately the same true discovery rate. However, ODT’s false discovery rate is lower. More instances help the ODT method to reduce its false discovery rate.
@@
Figure 14: Results on the synthetic tree data sets for increasing (a) number of training samples, and (b) number of numeric features; and (c) on the synthetic linear data sets for increasing number of training samples and numeric features, to evaluate Claim 3.
@@
Fig. 14 additionally shows that when learning from synthetic data with a linear separator as the ground truth, CART without a depth constraint achieves higher accuracy than ODT, but with more data availability, CART also generates much larger trees with only a small gain in accuracy. Increasing the number of features in the synthetic data makes the classification function harder to learn. The relative accuracy performance of the methods stays roughly the same, but CART requires many more nodes. In all cases, ODTs perform better than CART with the same depth constraint.
@@
Figure 16: Number of leaf nodes for CART for increasing training samples with and without tuning. These numbers far exceed the eight leaf nodes for ODTs in Fig. 15.
Fig. 15 shows the out-of-sample accuracies for increasing training sample sizes on three large real data sets. We train ODTs with a depth limit of three, and CART with and without a depth limit of three. When the methods are not tuned, the optimal approach overfits on small data sets, obtaining a lower test accuracy than depth-limited CART. However, with tuning, this effect disappears and the ODTs’ accuracy is consistently higher than CART’s (depth limited). The difference in performance between tuning and not tuning diminishes for larger training sets. Without a depth limit, CART continues to increase its accuracy for more data. More data does not help depth-limited CART since, at some point, the greedy decisions on what feature and threshold to split on do not change anymore. In those cases, the added data does not lead to different greedy decisions but only makes them more certain.
However, Fig. 16 also shows that CART continues to grow larger trees with up to tens of thousands of nodes. For the Electricity data set, for example, CART and ODTs have roughly similar accuracy for ten thousand training samples. However, the ODT has eight leaf nodes, whereas CART has over a hundred. These results contradict the observation by Oates and Jensen (1997) that greedy tree methods do not perform much better for more data but do yield larger trees with more data. For these data sets, we observe CART performing much better for more data while also resulting in larger trees.
In conclusion, these results refute Claim 3 that the difference between optimal methods and greedy diminished for more data. CART (without a depth limit) can improve performance over ODT with sufficient data. However, unconstrained CART uses unnecessary splits and can result in trees that are orders of magnitude larger than ODTs. Depth-constrained CART may fail to recover an accurate tree even with large training sets and the difference with ODTs does not diminish. Therefore, for both depth-constrained and unconstrained CART, we find that they remain different from ODTs with more data.
@@ 5.4.5 Model Complexity and Training Data
Figure 17: Testing Claim 4 on the synthetic data sets for ground truth trees of increasing complexity and training samples $n=50\cdot 2^{d}$, with $d$ the depth of the ground truth tree.
@@ 5.4.6 Overfitting
Figure 18: Testing Claim 5 on the synthetic tree data for increasing (a) feature noise, and (b) class noise; and (c) on the synthetic linear data for increasing feature and class noise.
We addressed overfitting before when observing the results in Fig. 15. These results showed that without hyperparameter tuning, ODTs are more prone to overfitting than greedy approaches when data is sparse. However, with tuning and with the same size constraint, we observe that ODTs perform better than greedy trees on average. We further analyze the risk of overfitting by comparing ODTs with greedy on the synthetic data with noise.
Fig. 18 shows how both approaches respond to increasing feature noise. For all amounts of feature noise, ODT obtains both a higher test accuracy and smaller trees than both CART approaches. The TDR is mostly similar, except for large amount of feature noise which causes the unconstrained CART to train larger trees which results in a higher TDR, but not in higher accuracy. ODT’s FDR is consistently better than CART’s.
Fig. 18 shows that for increasing amounts of class noise, ODT’s test accuracy is again consistently higher than both CART approaches. For large amounts of class noise, unconstrained CART obtains a lower test accuracy than both other approaches and also yields significantly larger trees.
For the synthetic linear data, Fig. 18 shows similar results. In both cases, with little noise, unconstrained CART achieves a higher accuracy but with a much larger tree. However, when either type of noise increases, ODT’s test accuracy becomes higher.
These results show that ODTs are not more sensitive to noise than greedy trees. Combined with the previous result on learning trees with a small training sample, this proves that Claim 5 is false when ODTs are properly tuned. Without tuning, given a fixed size limit, ODTs can overfit more than greedy trees. However, ODTs achieve the same performance as greedy trees at smaller size limits, which means there is no difference in overfitting after tuning. Therefore, the regularization strength of a fixed size on optimal and greedy trees is different and cannot be directly compared.
5.4.7 Question Length
@@
The question length of greedy trees remains (in practice) close to the optimal tree depth (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
Figure 19: The question length on the synthetic data is almost always the same, except for data sets with a strong noise component, mostly supporting Claim 6.
The question length is the expected number of branching nodes visited by a test instance. We evaluate Claim 6 about the question length of greedy trees on the synthetic data.
Fig. 19 shows that the question length of depth-constrained CART remains similar to the optimal approach. Also without a depth limit, CART’s question length in almost all cases –for varying training data size, number of features, and ground truth complexity– remains the same as the optimal approach. However, with much feature or class noise, CART’s question length is (much) larger than ODT’s. This difference is most pronounced for more than 30% class noise. Because Murthy and Salzberg (1995) tested with class noise up to 25% this effect was not previously observed.
Theoretically it can be shown that CART’s question length can be arbitrarily larger than optimal (Garey and Graham, 1974). But Claim 6 concerns practical use cases. Therefore, since the difference between CART and ODTs only arises for large noise, we consider Claim 6 supported by our experiments.
@@ 5.5 Scalability of Optimal Decision Trees
Since scalability is one of the limitations of ODTs, we test for what problem sizes the use of ODTs is practically feasible. Table 3 provides an overview of runtimes for the optimal method STreeD in terms of seconds, minutes, and hours when trained on synthetically generated data from a random decision tree. We find that training ODTs up to depth four remains practically feasible for data sets up to approximately 250 binary features for 100,000 instances and 150 binary features for one million instances.
We recollect the observation made before that ODTs are specifically good
  1. 1.
    for training small trees up to depth four (Claim 1);
  2. 2.
    for obtaining trees with a good accuracy-interpretability trade-off (Claim 2);
  3. 3.
    and increasingly so for data sets with many instances (Claim 3).
Therefore, we conclude that the best application of ODTs is finding small interpretable trees for medium to large data sets with relatively few but meaningful and well-prepared features.
@@
Table 3: Approximate magnitudes of runtimes for training ODTs using STreeD on synthetic data. s for (sub)seconds, m for minutes, and dashes for runtimes over two hours.
@@ 6 Conclusion
We experimentally evaluated many different approaches on how to train optimal decision trees (including several novel ones) and how they compare to greedy approaches. Based on the obtained insights, we provide recommendations for training and evaluating optimal and greedy methods (Recommendations 1-3).
@@
Since the design of greedy top-down induction (TDI) methods prevents direct accuracy optimization, the literature shows a large variety of learning objectives. We identified and analyzed nine such existing decision tree learning objectives and observed that the concavity of these objectives leads to counter-intuitive results. Based on this analysis, we evaluate two new non-concave objective functions. By optimizing ODTs for in-sample accuracy, we found significant improvements in out-of-sample accuracy over objectives such as entropy and Gini impurity. For noisy data sets, we show that objectives that have an additional regularizing effect (such as C4.5’s pessimistic binomial error) may improve out-of-sample performance, while also improving the sparsity of the trees.
@@
We further introduced two novel methods for tuning a tree’s complexity and evaluated these against four existing methods across 180 datasets with trees up to depth four, and 157 datasets with trees up to depth five. The experiments show that tuning has a statistically significant impact on the accuracy. The differences in accuracy between the six methods are small, but they may yield smaller or larger trees. Tuning only the depth is promising if runtime performance is a concern. Tests with synthetic noisy data shows that tuning the question length or the minimum support is significantly worse than the other approaches.
@@
Another important contribution of this work is establishing a set of best practices and the supporting analysis framework for comparing ODT and greedy methods. For this, we analyzed previous comparisons and then used the newly established best experimental practices (Recommendations 2) to test six claims from the literature, leading to the following insights. Our results confirm that (1) ODTs, on average, outperform greedy trees by 1-2% in out-of-sample accuracy for trees trained up to depth four. (2) ODTs, on average, have a better accuracy-interpretability trade-off, which we evaluate with a new metric: the size-weighted accuracy. We refute the claim that (3) differences between greedy and optimal diminish for more data. Depth-constrained greedy methods may fail to recover the true tree, even for large data sets, where ODTs succeed. Unlike previously claimed, (4) greedy methods do not maintain a similar accuracy when the data set size increases linearly with the true tree complexity. ODTs, on the other hand, do. When the complexity of a tree is properly tuned, we find no support for the claim that (5) ODTs are more prone to overfitting than greedy trees. Finally, we find supporting evidence for the claim that (6) the question length of greedy trees remains close to that of the optimal tree.
Although unrestricted greedy trees can outperform depth-limited ODTs in accuracy, they can quickly grow so large that they cannot be directly interpreted anymore. Random forests or neural networks already suffice if accuracy is the only concern. However, ODTs are an ideal candidate if interpretability is required, as they achieve a superior accuracy-interpretability trade-off over (unrestricted) greedy trees.